# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 2.2 From orthogonal polynomials to Lax Pairs
    $Y_{n}(z)=\begin{bmatrix}p_{n}(z)&\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{(w-z)2i\pi}\\ \widetilde{p}_{n-1}(z)&\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{(w-z)2i\pi}\end{bmatrix},$ (2.19)
    where $p_{n}(z)$ is the monic orthogonal polynomial of degree $n$ and $\widetilde{p}_{n-1}$ a polynomial of degree $\deg\widetilde{p}_{n-1}\leq n-1$; they are expressible explicitly in terms of the moments as follows
@@
    $\widetilde{p}_{n-1}(z)=\frac{-2i\pi}{\tau_{n}}\det\begin{bmatrix}\mu_{0}&\mu_{1}&\cdots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\cdots&\mu_{n}\\ \mu_{2}&\cdots&&\mu_{n+1}\\ \vdots\\ \mu_{n-2}&\cdots&&\mu_{2n-3}\\ 1&z&\cdots&z^{n-1}\end{bmatrix},$ (2.21)
@@
The matrix $Y_{n}(z;t)$ (we emphasize its dependence on $t$ as well) satisfies a first order ODE; indeed let us define
@@
We note that near $z=0,1,\infty$, the matrix $\Psi$ has the following expansion, which follows from the properties of $Y_{n}$ and the definition of $\theta$ (2.6)
$\Psi(z;t)=\left\{\begin{array}[]{lc}{\mathcal{O}}^{\times}(1)z^{\frac{{{K}}-{{\rho}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2z}\sigma_{3}},&z\to 0\\ {\mathcal{O}}^{\times}(1)(z-1)^{\frac{{{\rho}}}{2}\sigma_{3}},&z\to 1\\ \left(\mathbf{1}+\mathcal{O}(z^{-1})\right)z^{\left(n+\frac{{K}}{2}\right)\sigma_{3}},&z\to\infty\end{array}\right.$ (2.23)
@@
Proposition 2.1
The matrix $\Psi(z;t)$ in (2.22) satisfies the pair of first order PDEs
$\left\{\begin{array}[]{cc}\partial_{z}\Psi(z;t)=A(z;t)\Psi(z;t)\\ \partial_{t}\Psi(z;t)=B(z;t)\Psi(z;t)\end{array}\right.$ (2.24)
@@
$A(z;t)$ $:=-\frac{t}{z^{2}}G_{0}\sigma_{3}G_{0}^{-1}+G_{0}\left(\frac{{{K}}-{{\rho}}}{z}\sigma_{3}-\frac{t}{z}\Big{[}G_{0}^{-1}G_{1},\sigma_{3}\Big{]}\right)G_{0}^{-1}+\frac{{{K}}}{z-1}H_{0}\sigma_{3}H_{0}^{-1}+\frac{n+{{K}}}{z}\sigma_{3}$ (2.25)
$B(z;t)$ $:=\frac{1}{z}G_{0}\sigma_{3}G_{0}^{-1}.$ (2.26)
@@
Proof. It follows from the RHP 2.1 that $\Psi$ satisfies a boundary value problem
$\Psi(z_{+};t)=\Psi(z_{-};t)\begin{bmatrix}1&1\\ 0&1\end{bmatrix}.$ (2.27)
@@
Since the jump matrices are constant with respect to $z$, both $\Psi$ and $\partial_{z}\Psi$ satisfy the same jump relation and hence the matrix
$A(z;t):=\partial_{z}\Psi(z;t)\Psi^{-1}(z;t)$ (2.28)
extends analytically to an analytic function with at most isolated singularities at $z=0,1$. Thus, we have
$A(z;t)=Y_{n}^{\prime}Y_{n}^{-1}+Y_{n}\left(\frac{-t}{z^{2}}+\frac{{{\rho}}}{z(z-1)}+\frac{{{K}}}{z}\right)\sigma_{3}Y_{n}^{-1}.$ (2.29)
@@
This shows that the only singularities of $A(z)$ are a double pole at $z=0$ and a simple pole at $z=1$; near $z=\infty$, using the asymptotic behaviour of $Y_{n}$ as in (2.18), we conclude that
$A(z;t)=-\frac{t}{z^{2}}G_{0}\sigma_{3}G_{0}^{-1}+G_{0}\left(\frac{{{K}}-{{\rho}}}{z}\sigma_{3}-\frac{t}{z}\Big{[}G_{0}^{-1}G_{1},\sigma_{3}\Big{]}\right)G_{0}^{-1}+\frac{{{K}}}{z-1}H_{0}\sigma_{3}H_{0}^{-1}+\frac{n+{{K}}}{z}\sigma_{3}$ (2.30)
$G_{0}:=Y_{n}(0),\ \ \ G_{1}:=Y_{n}^{\prime}(0);\ \ \ \ H_{0}:=Y_{n}(1).$
A similar argument produces the expression for $B(z;t)$. $\blacksquare$
@@
In the pair of PDE’s (2.24) the $z$–equation is an ODE (considering $t$ as parameter) with an irregular singularity of Poincaré rank $2$ at $z=0$ and two Fuchsian singularities at $z=1,\infty$, respectively.
Comparison with the Japanese Lax Pair.
@@
In the work of the Japanese school [33] the $z$–component of the Lax pair for Painlevé V has also two Fuchsian and one second rank singularities, but with the positions reversed. More specifically, the Lax pair proposed in [33] is as follows [^4] We transcribe the results of [33] but we adapt their notation to our conventions. Note that the paper contains a couple of minor typographical errors: in (C.39) there should be an $x$ in front of the first term of $B$ and the sign of the $(1,2)$ entry of the next term should be the opposite. (see formulas (C.38–C.45) in loc. cit.)
$\partial_{\zeta}\Phi(\zeta;t)$ $=A_{{}_{JMU}}(\zeta;t)\Phi(\zeta;t)\ ,\ \ \qquad\partial_{t}\Phi(\zeta;t)=B_{{}_{JMU}}(\zeta;t)\Phi(\zeta;t)$
@@
$A_{{}_{JMU}}(\zeta;t)$ $=\frac{t}{2}\sigma_{3}+\frac{1}{\zeta}\begin{bmatrix}Z+\frac{\theta_{0}}{2}&-U(Z+\theta_{0})\\ \frac{Z}{U}&-Z-\frac{\theta_{0}}{2}\end{bmatrix}+\frac{1}{\zeta-1}\begin{bmatrix}-Z-\frac{\theta_{0}+\theta_{\infty}}{2}&UY\left(Z+\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\\ -\frac{Z+\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}}{UY}&Z+\frac{\theta_{0}+\theta_{\infty}}{2}\end{bmatrix}$
@@
$B_{{}_{JMU}}(\zeta;t)$ $=\frac{\zeta}{2}\sigma_{3}+\begin{bmatrix}0&-\frac{U}{t}\left(Z+\theta_{0}-Y\left(Z+\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\right)\\ \frac{1}{tUY}\left((Y-1)Z+\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right)&0\end{bmatrix},$ (2.31)
@@
where $Z=Z(t),\ Y=Y(t),\ U=U(t)$ satisfy a nonlinear first order system of ODEs in $t$ (C.40 in loc. cit.), which implies the fifth Painlevé equation for $Y$:
@@
$\frac{{\mathrm{d}}^{2}Y}{{\mathrm{d}}t^{2}}=\left(\frac{1}{2Y}+\frac{1}{Y-1}\right)\left(\frac{{\mathrm{d}}Y}{{\mathrm{d}}t}\right)^{2}-\frac{1}{t}\frac{{\mathrm{d}}Y}{{\mathrm{d}}t}+\frac{(Y-1)^{2}(\alpha Y+\frac{Y}{\beta})}{t^{2}}+\frac{\gamma Y}{t}+\frac{\delta Y(Y+1)}{Y-1}$
@@
$\alpha=\frac{1}{2}\left(\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)^{2};\ \beta=-\frac{1}{2}\left(\frac{\theta_{0}-\theta_{1}-\theta_{\infty}}{2}\right)^{2};\ \gamma=1-\theta_{0}-\theta_{1};\ \delta=-\frac{1}{2}.$ (2.32)
@@
$\Phi(\zeta;t)$ $=\mathcal{O}(1)\zeta^{\frac{\theta_{0}\sigma_{3}}{2}}\ ,$ $\zeta\to 0$
$\Phi(\zeta;t)$ $=\mathcal{O}(1)(\zeta-1)^{\frac{\theta_{1}\sigma_{3}}{2}}\ ,$ $\zeta\to 1$
$\Phi(\zeta;t)$ $=\left(\mathbf{1}+\frac{\Phi_{1}}{\zeta}+\mathcal{O}(\zeta^{-2})\right)\zeta^{-\frac{\theta_{\infty}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}\ ,$ $\zeta\to\infty.$ (2.33)
@@
and the equation admits the so–called sigma-form: indeed, introducing the new function
$\sigma(t)=tH_{V}-\frac{t}{2}(\theta_{0}+\theta_{\infty})+\frac{(\theta_{0}+\theta_{\infty})^{2}-\theta_{1}^{2}}{4},$ (2.35)
it can be verified that it satisfies ([33] formula (C.45))
$\left(t\frac{{\mathrm{d}}^{2}\sigma}{{\mathrm{d}}t^{2}}\right)^{2}=$ $\left(\sigma-t\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}+2\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}-(\theta_{\infty}+2\theta_{0})\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}+$
$-4\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\theta_{0}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right).$ (2.36)
In order to identify our Lax pair (2.24) with the Japanese one (2.2) it suffices to map $\zeta=\frac{1}{z}$ and suitably normalize our matrix $\Psi(z;t)$.
Lemma 2.1
@@
$\Phi(\zeta;t)=Y_{n}(0;t)^{-1}\Psi\left(\frac{1}{\zeta};t\right){\rm e}^{-i\frac{\pi}{2}{{\rho}}\sigma_{3}}$ (2.37)
transforms the Lax pair (2.24) into (2.2) with parameters
$\theta_{0}=-2n-{{K}};\qquad\theta_{1}={{\rho}};\qquad\theta_{\infty}={{K}}-{{\rho}}.$ (2.38)
This corresponds to the parameters $\alpha,\beta,\gamma$ in (2.2) as follows
$\alpha=\frac{(n+{{\rho}})^{2}}{2};\ \ \beta=-\frac{(n+{{K}})^{2}}{2};\ \ \gamma=1+2n+{{K}}-{{\rho}}.$ (2.39)
Proof. The map $\zeta=\frac{1}{z}$ maps $z=0$ to $\zeta=\infty$, $z=\infty$ to $\zeta=0$ and $z=1$ to $\zeta=1$. Thus the exponents of (formal) monodromy $\theta_{\{0,1,\infty\}}$ are the read off by matching the exponents in (2.23) and (2.33). $\blacksquare$
@@
Proposition 2.2
The Hankel determinant $\tau_{n}(t)$ in (2.9) is a polynomial tau function of the fifth Painlevé equation. In particular the Hamiltonian is
$H_{V}=-\frac{1}{2}{\rm tr}\left(Y_{n}^{-1}(0)Y^{\prime}_{n}(0)\sigma_{3}\right)+\frac{{\rho}}{2}=-\left(Y_{n}^{-1}(0)Y^{\prime}_{n}(0)\right)_{11}+\frac{{\rho}}{2}=\frac{{\mathrm{d}}}{{\mathrm{d}}t}\ln\tau_{n}(t)+\frac{{\rho}}{2}.$ (2.40)
Proof. We need to identify $\Phi_{1}$ in the expansion at $\zeta=\infty$ in (2.33) with the suitable expansion of $\Psi$ at $z=0$. Recalling the definition of $\Psi$ (2.22) we see that
$Y_{n}(0)^{-1}\Psi\left(\frac{1}{\zeta};t\right){\rm e}^{-\frac{i\pi}{2}{{\rho}}\sigma_{3}}=$ $Y_{n}(0)^{-1}\left(Y_{n}(0)+\frac{1}{\zeta}Y_{n}^{\prime}(0)+\mathcal{O}(\zeta^{-2})\right)\big{(}\zeta-1\big{)}^{\frac{{\rho}}{2}\sigma_{3}}\zeta^{-\frac{{K}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}=$
$=$ $\left(\mathbf{1}+\frac{Y_{n}(0)^{-1}Y_{n}^{\prime}(0)}{\zeta}+\mathcal{O}(\zeta^{-2})\right)\left(\mathbf{1}-\frac{{{\rho}}}{2\zeta}\sigma_{3}+\mathcal{O}(\zeta^{-2})\right)\zeta^{\frac{{{\rho}}-{{K}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}=$
$=$ $\left(\mathbf{1}+\frac{Y_{n}(0)^{-1}Y_{n}^{\prime}(0)-\frac{{\rho}}{2}\sigma_{3}}{\zeta}+\mathcal{O}(\zeta^{-2})\right)\zeta^{\frac{{{\rho}}-{{K}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}.$ (2.41)
Thus we conclude that
$\Phi_{1}=Y_{n}(0)^{-1}Y_{n}^{\prime}(0)-\frac{{\rho}}{2}\sigma_{3}.$ (2.42)
Then, according to (2.34), we must compute
$H_{{}_{V}}=-\frac{1}{2}{\rm tr}\left(\Phi_{1}\sigma_{3}\right)=-\frac{1}{2}{\rm tr}\left(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\sigma_{3}\right)+\frac{{\rho}}{2}=-\left(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\right)_{11}+\frac{{\rho}}{2},$ (2.43)
@@
Using that $\det Y_{n}\equiv 1$ and the formula (2.19) we deduce
@@
$(Y_{n}(0)^{-1}Y_{n}^{\prime}(0))_{11}=p^{\prime}_{n}(0)\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{w2i\pi}-{\widetilde{p}}\,^{\prime}_{\!n-1}(0)\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{w2i\pi}.$ (2.44)
The derivatives at $z=0$ of $p_{n},\widetilde{p}_{n-1}$ are obtained from their determinantal representations (2.20), (2.21):
$p^{\prime}_{n}(0)$ $={(-1)^{n+1}}\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}}$ (2.45)
@@
$\widetilde{p}^{\prime}_{n-1}(0)$ $=\dfrac{(-1)^{n+1}2\pi i}{\tau_{n}}\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}.$ (2.46)
@@
The contour integrals in (2.44) can be expressed as well in terms of determinants of the moments $\mu_{j}$ provided we extend their definition to all $j\in{\mathbb{Z}}$:

### v3

@@ 3. $(3)$
    $Y_{n}(z)=\begin{bmatrix}p_{n}(z)&\displaystyle\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{(w-z)2{\rm i}\pi}\vskip 2.84526pt\\ \widetilde{p}_{n-1}(z)&\displaystyle\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{(w-z)2{\rm i}\pi}\end{bmatrix}$, (2.7)
    where $p_{n}(z)$ is the monic orthogonal polynomial of degree $n$ and $\widetilde{p}_{n-1}$ a polynomial of degree $\deg\widetilde{p}_{n-1}\leq n-1$}. They are expressible explicitly in terms of the moments as follows:
@@
    $\widetilde{p}_{n-1}(z)=\frac{-2{\rm i}\pi}{\tau_{n}}\det\begin{bmatrix}\mu_{0}&\mu_{1}&\cdots&\mu_{n-1}\\ \mu_{1}&\mu_{2}&\cdots&\mu_{n}\\ \mu_{2}&\cdots&&\mu_{n+1}\\ \vdots\\ \mu_{n-2}&\cdots&&\mu_{2n-3}\\ 1&z&\cdots&z^{n-1}\end{bmatrix}$, (2.9)
@@
The matrix $Y_{n}(z;t)$ (we emphasize its dependence on $t$ as well) satisfies a first-order ODE. Indeed, let us define
@@
We note that near $z=0,1,\infty$, the matrix $\Psi$ has the following expansion, which follows from the properties of $Y_{n}$ and the definition of $\theta$ (2.2):
$\Psi(z;t)=\begin{cases}{\mathcal{O}}^{\times}(1)z^{\frac{{{K}}-{{\rho}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2z}\sigma_{3}},&z\to 0,\\ {\mathcal{O}}^{\times}(1)(z-1)^{\frac{{{\rho}}}{2}\sigma_{3}},&z\to 1,\\ \bigl(\mathbf{1}+\mathcal{O}\bigl(z^{-1}\bigr)\bigr)z^{(n+\frac{{K}}{2})\sigma_{3}},&z\to\infty,\end{cases}$ (2.11)
@@
Proposition 2.6.
The matrix $\Psi(z;t)$ in (2.10) satisfies the pair of first-order PDEs
$\partial_{z}\Psi(z;t)=A(z;t)\Psi(z;t),\qquad\partial_{t}\Psi(z;t)=B(z;t)\Psi(z;t),$ (2.12)
@@
$A(z;t):=-\frac{t}{z^{2}}G_{0}\sigma_{3}G_{0}^{-1}+G_{0}\left(\frac{{{K}}-{{\rho}}}{z}\sigma_{3}-\frac{t}{z}\big[G_{0}^{-1}G_{1},\sigma_{3}\big]\right)G_{0}^{-1}$
$\phantom{A(z;t):=}{}+\frac{{{K}}}{z-1}H_{0}\sigma_{3}H_{0}^{-1}+\frac{n+{{K}}}{z}\sigma_{3},$
$B(z;t):=\frac{1}{z}G_{0}\sigma_{3}G_{0}^{-1}.$
@@
It follows from the Riemann–Hilber Problem 2.4 that $\Psi$ satisfies a boundary value problem
$\Psi(z_{+};t)=\Psi(z_{-};t)\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$.
@@
Since the jump matrices are constant with respect to $z$, both $\Psi$ and $\partial_{z}\Psi$ satisfy the same jump relation and hence the matrix $A(z;t):=\partial_{z}\Psi(z;t)\Psi^{-1}(z;t)$ extends analytically to an analytic function with at most isolated singularities at $z=0,1$. Thus, we have
$A(z;t)=Y_{n}^{\prime}Y_{n}^{-1}+Y_{n}\left(\frac{-t}{z^{2}}+\frac{{{\rho}}}{z(z-1)}+\frac{{{K}}}{z}\right)\sigma_{3}Y_{n}^{-1}.$
@@
This shows that the only singularities of $A(z)$ are a double pole at $z=0$ and a simple pole at z=1{z=1}. Near $z=\infty$, using the asymptotic behaviour of $Y_{n}$ as in (2.6), we conclude that
$A(z;t)=-\frac{t}{z^{2}}G_{0}\sigma_{3}G_{0}^{-1}+G_{0}\left(\frac{{{K}}-{{\rho}}}{z}\sigma_{3}-\frac{t}{z}\Big[G_{0}^{-1}G_{1},\sigma_{3}\Big]\right)G_{0}^{-1}$
$\phantom{A(z;t)=}{}+\frac{{{K}}}{z-1}H_{0}\sigma_{3}H_{0}^{-1}+\frac{n+{{K}}}{z}\sigma_{3},$
$G_{0}:=Y_{n}(0),\qquad G_{1}:=Y_{n}^{\prime}(0),\qquad H_{0}:=Y_{n}(1).$
A similar argument produces the expression for $B(z;t)$. ∎
@@
In the pair of PDEs (2.12), the $z$-equation is an ODE (considering $t$ as parameter) with an irregular singularity of Poincaré rank $2$ at $z=0$ and two Fuchsian singularities at $z=1,\infty$, respectively.
@@
Comparison with the Japanese Lax pair. In the work of the Japanese school [33], the $z$-component of the Lax pair for Painlevé V has also two Fuchsian and one second rank singularities, but with the positions reversed. More specifically, the Lax pair proposed in [33] is as follows [^2] We transcribe the results of [33] but we adapt their notation to our conventions. Note that the paper contains a couple of minor typographical errors: in (C.39) there should be an $x$ in front of the first term of $B$ and the sign of the $(1,2)$ entry of the next term should be the opposite. (see formulas (C.38)–(C.45) in loc. cit.):
$\partial_{\zeta}\Phi(\zeta;t)=A_{{}_{\rm JMU}}(\zeta;t)\Phi(\zeta;t),\qquad\partial_{t}\Phi(\zeta;t)=B_{{}_{\rm JMU}}(\zeta;t)\Phi(\zeta;t),$
@@
$A_{{}_{\rm JMU}}(\zeta;t)=\frac{t}{2}\sigma_{3}+\frac{1}{\zeta}\begin{bmatrix}\displaystyle Z+\frac{\theta_{0}}{2}&-U(Z+\theta_{0})\vskip 2.84526pt\\ \displaystyle\frac{Z}{U}&\displaystyle-Z-\frac{\theta_{0}}{2}\end{bmatrix}$
[…]
$B_{{}_{\rm JMU}}(\zeta;t)=\frac{\zeta}{2}\sigma_{3}$ (2.13)
@@
$\quad{}+\begin{bmatrix}0&\displaystyle-\frac{U}{t}\left(Z+\theta_{0}-Y\left(Z+\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\right)\\ \displaystyle\frac{1}{tUY}\left((Y-1)Z+\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right)&0\end{bmatrix}$,
@@
where $Z=Z(t)$, $Y=Y(t)$, $U=U(t)$ satisfy a nonlinear first-order system of ODEs in $t$ ((C.40) in loc. cit.), which implies the fifth Painlevé equation for $Y$
@@
$\frac{{\mathrm{d}}^{2}Y}{{\mathrm{d}}t^{2}}=\left(\frac{1}{2Y}+\frac{1}{Y-1}\right)\left(\frac{{\mathrm{d}}Y}{{\mathrm{d}}t}\right)^{2}-\frac{1}{t}\frac{{\mathrm{d}}Y}{{\mathrm{d}}t}+\frac{(Y-1)^{2}\bigl(\alpha Y+\frac{Y}{\beta}\bigr)}{t^{2}}+\frac{\gamma Y}{t}+\frac{\delta Y(Y+1)}{Y-1},$
@@
$\alpha=\frac{1}{2}\left(\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)^{2},\qquad\beta=-\frac{1}{2}\left(\frac{\theta_{0}-\theta_{1}-\theta_{\infty}}{2}\right)^{2},$
$\gamma=1-\theta_{0}-\theta_{1},\qquad\delta=-\frac{1}{2}.$ (2.14)
@@
$\Phi(\zeta;t)=\mathcal{O}(1)\zeta^{\frac{\theta_{0}\sigma_{3}}{2}},\qquad\zeta\to 0,\qquad\Phi(\zeta;t)=\mathcal{O}(1)(\zeta-1)^{\frac{\theta_{1}\sigma_{3}}{2}},\qquad\zeta\to 1,$
$\Phi(\zeta;t)=\left(\mathbf{1}+\frac{\Phi_{1}}{\zeta}+\mathcal{O}\bigl(\zeta^{-2}\bigr)\right)\zeta^{-\frac{\theta_{\infty}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}},\qquad\zeta\to\infty.$ (2.15)
@@
and the equation admits the so-called sigma-form. Indeed, introducing the new function
$\sigma(t)=tH_{V}-\frac{t}{2}(\theta_{0}+\theta_{\infty})+\frac{(\theta_{0}+\theta_{\infty})^{2}-\theta_{1}^{2}}{4},$
it can be verified that it satisfies [33, formula (C.45)]
$\left(t\frac{{\mathrm{d}}^{2}\sigma}{{\mathrm{d}}t^{2}}\right)^{2}={}$ $\left(\sigma-t\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}+2\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}-(\theta_{\infty}+2\theta_{0})\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}$
$-4\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\theta_{0}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right).$
In order to identify our Lax pair (2.12) with the Japanese one (2.13), it suffices to map $\zeta=\frac{1}{z}$ and suitably normalize our matrix $\Psi(z;t)$.
Lemma 2.7.
@@
$\Phi(\zeta;t)=Y_{n}(0;t)^{-1}\Psi\left(\frac{1}{\zeta};t\right){\rm e}^{-{\rm i}\frac{\pi}{2}{{\rho}}\sigma_{3}}$
transforms the Lax pair (2.12) into (2.13) with parameters $\theta_{0}=-2n-{{K}}$, $\theta_{1}={{\rho}}$, $\theta_{\infty}={{K}}-{{\rho}}$. This corresponds to the parameters $\alpha$, $\beta$, $\gamma$ in (2.14) as follows:
$\alpha=\frac{(n+{{\rho}})^{2}}{2},\qquad\beta=-\frac{(n+{{K}})^{2}}{2},\qquad\gamma=1+2n+{{K}}-{{\rho}}.$
Proof.
The map $\zeta=\frac{1}{z}$ maps $z=0$ to $\zeta=\infty$, $z=\infty$ to $\zeta=0$ and $z=1$ to $\zeta=1$. Thus the exponents of (formal) monodromy $\theta_{\{0,1,\infty\}}$ are the read off by matching the exponents in (2.11) and (2.15). ∎
@@
Proposition 2.8.
The Hankel determinant $\tau_{n}(t)$ in (2.5) is a polynomial tau function of the fifth Painlevé equation. In particular, the Hamiltonian is
$H_{V}=-\frac{1}{2}{\rm tr}\left(Y_{n}^{-1}(0)Y^{\prime}_{n}(0)\sigma_{3}\right)+\frac{{\rho}}{2}=-\left(Y_{n}^{-1}(0)Y^{\prime}_{n}(0)\right)_{11}+\frac{{\rho}}{2}=\frac{{\mathrm{d}}}{{\mathrm{d}}t}\ln\tau_{n}(t)+\frac{{\rho}}{2}.$
Proof.
We need to identify $\Phi_{1}$ in the expansion at $\zeta=\infty$ in (2.15) with the suitable expansion of $\Psi$ at $z=0$. Recalling the definition of $\Psi$ (2.10), we see that
$Y_{n}(0)^{-1}\Psi\left(\frac{1}{\zeta};t\right){\rm e}^{-\frac{{\rm i}\pi}{2}{{\rho}}\sigma_{3}}$
$\qquad=Y_{n}(0)^{-1}\left(Y_{n}(0)+\frac{1}{\zeta}Y_{n}^{\prime}(0)+\mathcal{O}\bigl(\zeta^{-2}\bigr)\right)(\zeta-1)^{\frac{{\rho}}{2}\sigma_{3}}\zeta^{-\frac{{K}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}$
$\qquad=\left(\mathbf{1}+\frac{Y_{n}(0)^{-1}Y_{n}^{\prime}(0)}{\zeta}+\mathcal{O}\bigl(\zeta^{-2}\bigr)\right)\left(\mathbf{1}-\frac{{{\rho}}}{2\zeta}\sigma_{3}+\mathcal{O}\bigl(\zeta^{-2}\bigr)\right)\zeta^{\frac{{{\rho}}-{{K}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}$
$\qquad=\left(\mathbf{1}+\frac{Y_{n}(0)^{-1}Y_{n}^{\prime}(0)-\frac{{\rho}}{2}\sigma_{3}}{\zeta}+\mathcal{O}\bigl(\zeta^{-2}\bigr)\right)\zeta^{\frac{{{\rho}}-{{K}}}{2}\sigma_{3}}{\rm e}^{\frac{t}{2}\zeta\sigma_{3}}.$
Thus we conclude that $\Phi_{1}=Y_{n}(0)^{-1}Y_{n}^{\prime}(0)-\frac{{\rho}}{2}\sigma_{3}$. Then, according to (2.16), we must compute
$H_{{}_{V}}=-\frac{1}{2}{\rm tr}\left(\Phi_{1}\sigma_{3}\right)=-\frac{1}{2}{\rm tr}\left(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\sigma_{3}\right)+\frac{{\rho}}{2}=-\left(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\right)_{11}+\frac{{\rho}}{2},$
@@
Computation of $\boldsymbol{\bigl(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\bigr)_{11}}$. Using that $\det Y_{n}\equiv 1$ and the formula (2.7), we deduce
@@
$\bigl(Y_{n}(0)^{-1}Y_{n}^{\prime}(0)\bigr)_{11}=p^{\prime}_{n}(0)\oint_{\gamma}\frac{\widetilde{p}_{n-1}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{w2{\rm i}\pi}-{\widetilde{p}}^{\prime}_{\!n-1}(0)\oint_{\gamma}\frac{p_{n}(w){\rm e}^{\theta(w)}{\mathrm{d}}w}{w2{\rm i}\pi}.$ (2.17)
The derivatives at $z=0$ of $p_{n}$, $\widetilde{p}_{n-1}$ are obtained from their determinantal representations (2.8), (2.9)
$p^{\prime}_{n}(0)={(-1)^{n+1}}\dfrac{\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n+1}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-1}&\mu_{n+1}&\dots&\mu_{2n-1}\end{vmatrix}}{\tau_{n}},$
@@
$\widetilde{p}^{\prime}_{n-1}(0)=\dfrac{(-1)^{n+1}2\pi{\rm i}}{\tau_{n}}\begin{vmatrix}\mu_{0}&\mu_{2}&\dots&\mu_{n-1}\\ \mu_{1}&\mu_{3}&\dots&\mu_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \mu_{n-2}&\mu_{n}&\dots&\mu_{2n-3}\end{vmatrix}$.
@@
The contour integrals in (2.17) can be expressed as well in terms of determinants of the moments $\mu_{j}$ provided we extend their definition to all $j\in{\mathbb{Z}}$